OpenStudy (anonymous):

Write the inverse function for the function, ƒ(x) =1/2 x + 4. Then, find the value of ƒ^-1(4)

OpenStudy (amorfide):

to find the inverse of a function f(x)=x+3 we let f(x)=y then we want to make x the subject y=x+3 y-3=x now we replace x with f^-1(x) and replace y with x f^-1(x)=x-3

OpenStudy (amorfide):

so do the same for your question and let me know what you get

OpenStudy (amorfide):

\[f(x)=\frac{ 1 }{ 2 }x + 4\] ley y=f(x) \[y=\frac{ 1 }{ 2 }x + 4\] make x the subject

OpenStudy (anonymous):

so x= 1/2 y +4

OpenStudy (amorfide):

no

OpenStudy (amorfide):

do you know how to rearrange equations?

OpenStudy (anonymous):

not really and the notes i have don't really give me anything to work with

OpenStudy (amorfide):

you need to understand that addition and subtraction are opposite multiply and divide are opposites

OpenStudy (amorfide):

so if you have

OpenStudy (amorfide):

y=x+3

OpenStudy (amorfide):

make x the subject

OpenStudy (amorfide):

you must subtract 3 on both sides

OpenStudy (amorfide):

y-3=x

OpenStudy (amorfide):

if you have y=2x+4

OpenStudy (amorfide):

we subtract 4 on both sides, y-4=2x now we have 2 lots of x we want x on its own, so we must divide by 2

OpenStudy (amorfide):

\[\frac{ 1 }{ 2 }(y-4)=x\]

OpenStudy (amorfide):

so in your question

OpenStudy (amorfide):

\[y=\frac{ 1 }{ 2 }x+4\]

OpenStudy (amorfide):

can you make x the subject

OpenStudy (anonymous):

so it would start with subtracting 4 y-4= 1/2 x then divide 1/2 (y-4)/2=x Am I close?

OpenStudy (amorfide):

if you divide by 2 you are correct you said divide by 1/2, which is equivalent to multiplying by 2 so you actually have 2(y-4)=x

OpenStudy (amorfide):

so now that we have 2(y-4)=x

OpenStudy (amorfide):

we replace x with f^-1(x) and replace y with x

OpenStudy (amorfide):

2(x-4)=f^-1(x) this is our inverse function

OpenStudy (amorfide):

so in the question it wanted you to find f^-1(4) in this function so substitute x=4

OpenStudy (anonymous):

ok so the function would look like f^-1(4)=2(4-4) then you would multiply the 2 into the parentheses, right?

OpenStudy (amorfide):

well what is 4-4

OpenStudy (anonymous):

0 so f^-1(4)=0

OpenStudy (amorfide):

yes well done! thank you for testimony but please can you give me a medal for best answer?

OpenStudy (amorfide):

if you need more practice with inverse functions I can help you out, just let me know

OpenStudy (amorfide):

just a little guide if you have a function f(x)=2x+4 think of it as, you start with a number you multiply it by 2 then you add 4 so your starting number is x you multiply it by 2 then you added 4 this gives you a new number which is represented as f(x) so the inverse function gets you back to the original number so you do the opposite the opposite of add 4 is subtract 4 the opposite of multiply by 2 is divide by 2 so we have \[f^{-1}(x)=\frac{ 1 }{ 2 }(x-4)\] it is easier if you turn your given function into a word problem so you can see what you have to do to get back to the original number which would be your inverse function hope this helps a little more